# Critical Sections Are Not Per-Thread: A Trace Semantics for Lock-Based Concurrency

### 13 Mar 2026 v1

@@ 1. Introduction
For example, consider the program in Figure 1. We assume that threads are numbered where thread 1 is the thread in which the main function executes. At line 18 another thread 2 is created. Thread 2 acquires lock m1 and then acquires lock m2 before releasing the locks. We can argue that the lock acquire operation at line 6 is part of the critical section with entry point at line 5 and exit point at line 8.
Entry/exit points of critical sections are pairs of lock/unlock operations that operate on the same lock variable and belong to the same thread. Anything that must occur between the entry and exit points in all execution traces belongs to the critical section. The common assumption is that ‘must inbetween’ coincides with thread order, and therefore that critical sections are per-thread. This assumption is implicit and fundamental to the standard lock set construction [DS91].
@@ 2. Events and Traces
The projection of $T$ onto thread $\tau$, denoted $\mathit{proj}(T,\tau)$, restricts $T$ to events $e$ with where $e=(\tau,op)$. We say trace $T$ prefixes trace $T^{\prime}$ if $T^{\prime}=T\mathbin{\cdot}T^{\prime\prime}$ for some trace $T^{\prime\prime}$. We define $\mathit{thds}(T)=\{\tau\mid\exists e=(\tau,op)\in T\}$. For example, $\mathit{proj}(T_{3},\tau_{2})=[e_{2}]$ and $\mathit{evts}(T_{3})=\{e_{1},e_{6},e_{7},e_{8},e_{9},e_{10},e_{11},e_{2}\}$ and $\mathit{thds}(T_{1})=\{\tau_{1},\tau_{2},\tau_{3}\}$.
@@
For every $\tau\in\mathit{thds}(T^{\prime})$, $\mathit{proj}(T^{\prime},\tau)$ prefixes $\mathit{proj}(T,\tau)$.
@@ 3. Critical Section and Lock Set
We consider a specific program run represented by some well-formed trace $T$. In trace $T$ we wish to identify critical sections that protect events. The entry and exit points of a critical section are easily identified by matching lock and unlock events.
@@
[Matching Lock and Unlock Events] Let $T$ be a well-formed trace with events $l=(t,\mathtt{lock}(m)),u=(t,\mathtt{unlock}(m))\in T$. We say that $l$ and $u$ are matching lock and unlock events, written $\mathit{match}(l,u)_{T}$, if $l<_{tr}^{T}u$ and there exists no $u^{\prime}=(t,\mathtt{unlock}(m))\in T$ such that $l<_{tr}^{T}u^{\prime}<_{tr}^{T}u$. Due to well-formedness, each lock event has its distinct matching unlock event and vice versa. To avoid special treatment of ’open’ critical sections where the unlock event is missing we impose the following additional condition on traces:
WF-Acq2:
For all $l=(t,\mathtt{lock}(m))\in T$ there exists $u=(t,\mathtt{unlock}(m))\in T$ where $l<_{tr}^{T}u$.
The extra condition WF-Acq2 imposes no restriction as we can always add missing unlock events at the end of a trace.
Informally, an event is part of a critical section if it is enclosed by a lock event and its matching unlock event. In such a situation we say that the event holds the lock. The set of locks held is commonly referred to as the lock set. We first review the existing per-thread definition of a critical section and the resulting lock set construction and then introduce our more general trace-based characterization.
@@ 3.1. Per-Thread Definition
[Per-Thread Critical Section] Let $T$ be a well-formed trace with events $l=(\tau,\mathtt{lock}(m)),u=(\tau,\mathtt{unlock}(m))\in T$ where $\mathit{match}(l,u)_{T}$. We say that $e\in T$ is in the per-thread critical section guarded by $l$ and $u$ in thread $\tau$, written $e\in CS_{T}^{\tau}(l,u)$, if $l<_{tr}^{T}e<_{tr}^{T}u$ and $e=(\alpha,\tau,op)$ for some operation in the same thread as $l$ and $u$.
@@
[Per-Thread Locks Held] Let $T$ be a well-formed trace with an event $e=(\tau,op)\in T$. The per-thread lock set for an event $e\in T$ in thread $\tau$ is defined by $LH_{T}^{\tau}(e)=\{m\mid\exists l=(\tau,\mathtt{lock}(m)),u=(\tau,\mathtt{unlock}(m))\in T.e\in CS_{T}^{\tau}(l,u)\}$.
The superscript $\tau$ in $CS_{T}^{\tau}(l,u)$ and $LH_{T}^{\tau}(e)$ highlights that the above definitions are restricted to events that are in the same thread $\tau$.
@@
We can show that the lock set is stable in case program executions follows a different schedule of events.
Lemma 1.
Let $T$ be a well-formed trace with some event $e=(\tau,op)\in T$ where $m\in LH_{T}^{\tau}(e)$. Then, $m\in LH_{T^{\prime}}^{\tau}(e)$ for any reordering $T^{\prime}\in\mathit{crp}(T)$ where $e\in T^{\prime}$ and $T^{\prime}$ additionally satisfies WF-Acq2.
Proof 3.1.
By assumption, there exists $l=(\tau,\mathtt{lock}(m)),u=(\tau,\mathtt{unlock}(m))\in T$ such that $\mathit{match}(l,u)_{T}$ and $l<_{tr}^{T}e<_{tr}^{T}u$. Each $T^{\prime}\in\mathit{crp}(T)$ maintains the program order condition CRP-PO and with the additional condition WF-Acq2 we find that $l<_{tr}^{T^{\prime}}e<_{tr}^{T^{\prime}}u$. Hence, $e\in CS_{T^{\prime}}^{\tau}(l,u)$ and thus $m\in LH_{T^{\prime}}^{\tau}(e)$.
However, the per-thread lock set definition is semantically incomplete. For the running example in Figure 2 we find that for all reorderings $T^{\prime}\in\mathit{crp}(T)$ that additionally satisfy WF-Acq2 we have that event $e_{8}$ is surrounded by events $e_{6}$ and $e_{11}$. That is, $e_{6}<_{tr}^{T^{\prime}}e_{8}<_{tr}^{T^{\prime}}e_{11}$. This means that based on the observed execution represented by trace $T_{1}$, event $e_{8}$ is always protected by lock $m_{1}$ not matter which alternative schedule of events is taken. Due to the per-thread restriction of the lock set $LH_{T_{1}}^{\tau_{3}}(e_{8})$ lock $m_{1}$ is not an element in this set.
@@ 3.2. Trace-Based Characterization
[Critical Section] Let $T$ be a well-formed trace with events $l=(\tau,\mathtt{lock}(m)),u=(\tau,\mathtt{unlock}(m))\in T$ where $\mathit{match}(l,u)_{T}$. We say that $e\in T$ is in the critical section guarded by $l$ and $u$, written $e\in CS_{T}(l,u)$, if $l<_{crp}^{T}e<_{crp}^{T}u$.
@@
[Locks Held] Let $T$ be a well-formed trace with an event $e=(\tau,op)\in T$. The lock set for an event $e\in T$ in thread $\tau$ is defined by $LH_{T}(e)=\{m\mid\exists l=(\tau,\mathtt{lock}(m)),u=(\tau,\mathtt{unlock}(m))\in T.e\in CS_{T}(l,u)\}$.
Lemma 2.
Let $T$ be a well-formed trace with an event $e=(\tau,op)\in T$. Then, $LH_{T}^{\tau}(e)\subseteq LH_{T}(e)$.
The subset relation is strict as shown by the following example.
{exa}
For trace $T_{1}$ in Figure 2 and events $e_{3}=(\tau_{2},\mathtt{lock}(m_{2})),e_{8}=(\tau_{3},\mathtt{lock}(m_{2}))\in T_{1}$ we find that $LH_{T_{1}}(e_{3})=\{m_{1}\}$ and $LH_{T_{1}}(e_{8})=\{m_{1}\}$.
Thanks to the lifted per-thread restriction, lock set is now semantically complete. We say that lock $m$ protects event $e\in T$ if there exist $l=(\tau,\mathtt{lock}(m)),u=(\tau,\mathtt{unlock}(m))\in T$ where $\mathit{match}(l,u)_{T}$ and for all $T^{\prime}\in\mathit{crp}(T)$ where $l,u\in T^{\prime}$ we have that $l<_{crp}^{T^{\prime}}e<_{crp}^{T^{\prime}}u$. Any lock $m$ that protects some event is in the event’s lock set. The following result follows by definition.
Lemma 3 (Semantic Completeness of Trace-Based Lock Set).
Let $T$ be a well-formed trace and $e\in T$. If lock $m$ protects $e$ then $m\in LH_{T}(e)$.
We can also restate lock set stability.
Lemma 4 (Lock Set Stability).
Let $T$ be a well-formed trace with some event $e=(\tau,op)\in T$ where $m\in LH_{T}(e)$. Then, $m\in LH_{T^{\prime}}(e)$ for any reordering $T^{\prime}\in\mathit{crp}(T)$ where $e\in T^{\prime}$ and $T^{\prime}$ additionally satisfies WF-Acq2.

### 20 Mar 2026 v2

@@ 1. Introduction
For example, consider the program in Figure 1. We assume that threads are numbered where thread 1 is the thread in which the main function executes. At line 18 another thread 2 is created. Thread 2 acquires lock m1 and then acquires lock m2 before releasing the locks. The lock acquire operation at line 6 is part of the critical section with entry point at line 5 and exit point at line 8.
Entry and exit points of critical sections are pairs of lock and unlock operations on the same lock variable that belong to the same thread. Any event that must occur between the entry and exit points in all execution traces belongs to the critical section. The common assumption is that ‘must occur between’ coincides with thread order, and therefore that critical sections are per-thread. This assumption is implicit and fundamental to the standard lock set construction [DS91], yet it is rarely stated explicitly in the literature.
@@ 2. Events and Traces
The projection of $T$ onto thread $\tau$, denoted $\mathit{proj}_{T}(\tau)$, restricts $T$ to events $e$ where $e=(\tau,op)$. We say trace $T$ prefixes trace $T^{\prime}$ if $T^{\prime}=T\mathbin{\cdot}T^{\prime\prime}$ for some trace $T^{\prime\prime}$. We define $\mathit{thds}(T)=\{\tau\mid\exists e=(\tau,op)\in T\}$. For example, $\mathit{proj}_{T_{3}}(\tau_{2})=[e_{2}]$ and $\mathit{evts}(T_{3})=\{e_{1},e_{6},e_{7},e_{8},e_{9},e_{10},e_{11},e_{2}\}$ and $\mathit{thds}(T_{1})=\{\tau_{1},\tau_{2},\tau_{3}\}$.
@@
For every $\tau\in\mathit{thds}(T^{\prime})$, $\mathit{proj}_{T^{\prime}}(\tau)$ prefixes $\mathit{proj}_{T}(\tau)$.
@@ 3. Critical Section and Lock Set
We consider a specific program run represented by some well-formed trace $T$. In trace $T$ we wish to identify critical sections that protect events. The entry and exit points of a critical section are identified by matching lock and unlock events $l$ and $u$ where $l$ and $u$ operate on the same lock and are nearest to each other. By well-formedness events $l$ and $u$ must belong to the same thread. In case the program exited prematurely, a lock event $u$ might not have a matching unlock event $l$. In such a case, we consider the last event of the thread the lock event $u$ is in as the exit point.
@@
[Entry and Exit Points] Let $T$ be a well-formed trace with lock event $l=(\tau,\mathtt{lock}(m))\in T$. We refer to $l$ as the entry point of a critical section on lock $m$. The exit point of $l$ in trace $T$ is obtained via function $\mathit{exit}_{T}(l)$ defined below.
[…]
We assume that $\mathit{exit}_{T}()$ and $\mathit{find}_{T}()$ describe a family of functions indexed by some trace $T$. Function $\mathit{exit}_{T}(l)$ scans the trace for entry point $l$ and then returns the exit point via helper function $\mathit{find}_{T}(\tau,m)$.
{exa}
For trace $T_{1}$ from Figure 2 we find that $\mathit{exit}_{T_{1}}(e_{6})=e_{11}$, $\mathit{exit}_{T_{1}}(e_{8})=e_{9}$, $\mathit{exit}_{T_{1}}(e_{2})=e_{5}$ and $\mathit{exit}_{T_{1}}(e_{3})=e_{4}$. For trace $T_{3}=[e_{1},e_{6},e_{7},e_{8},e_{9},e_{10},e_{11},e_{2}]$ from Example 2 we find that $\mathit{exit}_{T_{3}}(e_{6})=e_{11}$ and $\mathit{exit}_{T_{3}}(e_{8})=e_{9}$ and $\mathit{exit}_{T_{3}}(e_{2})=e_{2}$.
The exit point for each entry point is stable under trace reorderings.
Lemma 1 (Stability of Exit Points).
Let $T$ be a well-formed trace with lock event $l=(t,\mathtt{lock}(m))\in T$. For any reordering $T^{\prime}\in\mathit{crp}(T)$ where $\mathit{exit}_{T}(l)\in T^{\prime}$ we have that $\mathit{exit}_{T}(l)=\mathit{exit}_{T^{\prime}}(l)$.
Proof 3.1.
By construction we have that $\mathit{exit}_{T}(l)=\mathit{exit}_{\mathit{proj}_{T}(\tau)}(l)$. For any $T^{\prime}\in\mathit{crp}(T)$ we have that $\mathit{proj}_{T^{\prime}}(\tau)$ is a prefix of $\mathit{proj}_{T}(\tau)$. By assumption $\mathit{exit}_{T}(l)\in T^{\prime}$ and then we can immediately follow that $\mathit{exit}_{T}(l)=\mathit{exit}_{T^{\prime}}(l)$.
We can argue that critical sections are uniquely identified by their entry and exit points. An event is part of a critical section if the event is enclosed by the entry and its corresponding exit point. In such a situation we say that the event is protected by the lock that is acquired at the entry point.
{defi}
[Lock Protection] Let $T$ be a well-formed trace with event $e\in T$. We say that $e$ is protected by lock $m$ if there exists lock event $l=(\tau,\mathtt{lock}(m))\in T$ and one of the following cases applies: (LP1) for all $T^{\prime}\in\mathit{crp}(T)$ where $\mathit{exit}_{T}(l)\in T^{\prime}$ we have that $l<_{tr}^{T^{\prime}}e<_{tr}^{T^{\prime}}\mathit{exit}_{T}(l)$, or (LP2) $e=\mathit{exit}_{T}(l)$ and $e\neq(\tau,\mathtt{unlock}(m))$. Case (LP2) covers the case that a lock event $l$ does not have a matching unlock event. Then, $\mathit{exit}_{T}(l)$ equals the last event $\mathit{exit}_{T}(l)$ in $\tau$ and acts as a delimiter for the critical section. In such a case, we consider $\mathit{exit}_{T}(l)$ part of the critical section and protected by lock $m$.
The set of locks that protect an event is commonly referred to as the lock set. Next, we review the existing per-thread definition of a critical section and the resulting lock set construction where the check for enclosed events only considers events that are in the same thread as the exit point. We point out the semantic incompleteness of this assumption and introduce our more general trace-based characterization for which we can establish semantic completeness.
@@ 3.1. Per-Thread Definition
[Per-Thread Critical Section] Let $T$ be a well-formed trace with lock event $l=(\tau,\mathtt{lock}(m))\in T$. We say that $e\in T$ is in the per-thread critical section guarded by $l$ in thread $\tau$, written $e\in CS_{T}^{\tau}(l)$, if one of the following cases applies: (PCS1) $\mathit{proj}_{T}(\tau)=T_{1}\mathbin{\cdot}[l]\mathbin{\cdot}T_{2}\mathbin{\cdot}[\mathit{exit}_{T}(l)]\mathbin{\cdot}T_{3}$ for some traces $T_{1}$, $T_{2}$ and $T_{3}$ where $e\in T_{2}$, or (PCS2) $e=\mathit{exit}_{T}(l)$ and $e\neq(\tau,\mathtt{unlock}(m))$. To deal with ’open’ critical sections where the matching unlock event is missing, we use two cases similar to the prior Definition 3.1.
@@
[Per-Thread Locks Held] Let $T$ be a well-formed trace with an event $e=(\tau,op)\in T$. The per-thread lock set for an event $e\in T$ in thread $\tau$ is defined by $LH_{T}^{\tau}(e)=\{m\mid\exists l=(\tau,\mathtt{lock}(m))\in T.e\in CS_{T}^{\tau}(l)\}$.
The superscript $\tau$ in $CS_{T}^{\tau}(l)$ and $LH_{T}^{\tau}(e)$ highlights that the above definitions are restricted to events that are in the same thread $\tau$.
@@
Each lock in the per-thread lock set protects the associated event as shown by the following result.
Lemma 2 (Semantic Correctness of Per-Thread Lock Set).
Let $T$ be a well-formed trace with some event $e=(\tau,op)\in T$ where $m\in LH_{T}^{\tau}(e)$. Then, $e$ is protected by lock $m$.
Proof 3.2.
By assumption, there exists $l=(\tau,\mathtt{lock}(m))\in T$ such that either (PCS1) $\mathit{proj}_{T}(\tau)=T_{1}\mathbin{\cdot}[l]\mathbin{\cdot}T_{2}\mathbin{\cdot}[\mathit{exit}_{T}(l)]\mathbin{\cdot}T_{3}$ for some traces $T_{1}$, $T_{2}$ and $T_{3}$ where $e\in T_{2}$, or (PCS2) $e=\mathit{exit}_{T}(l)$ and $e\neq(\tau,\mathtt{unlock}(m))$.
Consider $T^{\prime}\in\mathit{crp}(T)$ where $\mathit{exit}_{T}(l)\in T^{\prime}$. Each $T^{\prime}\in\mathit{crp}(T)$ maintains the program order condition CRP-PO. Hence, $\mathit{proj}_{T^{\prime}}(\tau)=T_{1}\mathbin{\cdot}[l]\mathbin{\cdot}T_{2}[\mathit{exit}_{T}(l)]\mathbin{\cdot}T_{3}$. Immediately, we can conclude that case (PCS1) implies (LP1) and (PCS2) implies (LP2). Thus, $e$ is protected by lock $m$.
However, the per-thread lock set definition is semantically incomplete.
{exa}
Consider trace $T_{1}$ in Figure 2 where $e_{11}=\mathit{exit}_{T_{1}}(e_{6})$. Due to the fork-join dependency we find that for all reorderings $T^{\prime}\in\mathit{crp}(T_{1})$ where $e_{11}\in T^{\prime}$ we have that event $e_{8}$ is surrounded by events $e_{6}$ and $e_{11}$. That is, $e_{6}<_{tr}^{T^{\prime}}e_{8}<_{tr}^{T^{\prime}}e_{11}$. This means that $e_{8}$ is protected by lock $m$, however, $m\not\in LH_{T_{1}}^{e_{8}}($).
@@ 3.2. Trace-Based Characterization
[Critical Section] Let $T$ be a well-formed trace with lock event $l=(\tau,\mathtt{lock}(m))\in T$. We say that $e\in T$ is in the critical section guarded by $l$, written $e\in CS_{T}(l)$, if one of the following two cases applies: (CS1) $l<_{crp}^{T}e<_{crp}^{T}\mathit{exit}_{T}(l)$, or (CS2) $e=\mathit{exit}_{T}(l)$ and $e\neq(\tau,\mathtt{unlock}(m))$.
@@
[Locks Held] Let $T$ be a well-formed trace with an event $e=(\tau,op)\in T$. The lock set for an event $e\in T$ in thread $\tau$ is defined by $LH_{T}(e)=\{m\mid\exists l=(\tau,\mathtt{lock}(m))\in T.e\in CS_{T}(l)\}$.
Lemma 3 (Semantic Correctness and Completeness of Lock Set).
Let $T$ be a well-formed trace with some event $e=(\tau,op)\in T$. Then, $m\in LH_{T}(\tau){e}$ iff $e$ is protected by lock $m$.
Proof 3.3.
We observe that case (LP1) in Definition 3.1 is equivalent to case (CS1) in Definition 3.2. The same applies to (LP2) and (CS2). Hence, $m\in LH_{T}^{\tau}(e)$ iff $e$ is protected by lock $m$.
An immediate consequence is that the per-thread lock set construction is a strict subset of the trace-based lock set construction.
From Lemma 2 and Lemma 3 we derive that $LH_{T}^{\tau}(e)\subseteq LH_{T}(e)$ for any $e\in T$. The subset relation is strict as shown by Example 3.1.